value of each years coupon payiments b. The yield to maturity is a measure of a bonds total return, only including the coupon income. . If you buy the bond today and hold it to maturity, your return will be yield to d. The relationship between price and yield is that the higher the price you pay for a bond, the higher the yield 1o. Which one of the following statements is correet regarding interest rates and bond values? a. When the interest rate rises, the present value of the payments to be recetved by the bondholder falls and bond prices fall. A decline in the interest rate decreases the present value of those payments and results in a higher price Any change in interest rate has a greater impact on the price of long-term bonds than the price of short-term bonds A change in interest rates has only a modest impact on the present value of long-term cash flows b. c. d. 11. Which of the following is not a element of a good corporate governance? a. Legal requirements b. Employee motivation. c. Boards of directors. d. Activist shareholders 12. A share of common stock paid a dividend of $4 at the end of last year. Dividends are expected to grow at a constant rate of 8% per year indefinitely. If investors require a rate of return of 10% what is the current share price? a. $40 b. $50 c. $200 d. $216 13. Which of the following statements about the intrinsic value of the share is correct? a. Intrinsic value is just the present value of the dividend payments anticipated by the investor in the stock b. In competitive market, the perfect price of the stock will equal to its intrinsic c. In competitive market, the perfect price of the stock will higher than its d. In competitive market, the perfect price of the stock will lower than its value. intrinsic value intrinsic value. 14. Which of the following statements is not correct regarding the P/E ratio? a. The P/E ratio is used to calculate the number of times the price being paid for 310

MULTIPLE CHOICES
Answer #1

(10): The answer is option “a” – when the interest rate rises the present value of the payments to be received by the bondholder falls and bond prices fall.

Explanation: Changes in interest rates affect the present value of the coupon payments but not the payment themselves. There exists an inverse relationship between interest rates and bond prices.

(11): The answer is option “b” – employee motivation.

Explanation: This is because employee motivation is not a part of accountability and supervision. Elements of corporate governance are good board practices, well-defined shareholders rights, board commitment, transparent disclosure etc.

(12): The answer is option “d” - $216

Explanation: D1 = $4*(1+8%) = 4.32

Thus current price = D1/r-g = 4.32/10%-8% = 4.32/2% = $216

(13): The answer is option “a” – Intrinsic value is just the present value of the dividend payments anticipated by the investor in the stock.

Explanation: This is as per the dividend discount model and as per this model the intrinsic value of the stock = sum of all the present values of future expected dividends.